1. Solve the following Differential Equations.

a) y - y’ – 2y = 5e2x b) y +16 y = 4 cos x c) y – 4y+3y=9x² +4, y(0) =6, y(0)=8

2. Use the variation of parameters method to find the general solution to the given differential equation.

y + y = tan?(x)

3.

Determine the general solution to the system x = Ax for the given matrix A. -1 2 А 2 2
Answer #1

page i 2.2 @ yl-y !-2y = 5 e 2x auxiliarly ean to solve homo. diff equation yll - 91-2y zo is m2 - m -2=0 m²-2m+ m -2 to mcm(Nagpur Universit: S- ons. Method of Variation of Par cosec page2 = clex teze2x + 5 e 2x (b = 4 cosx yu+t6y auxiliary egn forраа ез Оy - 4узу 9х2 4 со саlts) = 8 auxiliarely ean is m2 - 4 m +8 EO rm 2 - 3 m - m + 3 = D rm C -3) - 1 (от -3) = 0 Cm -Mersin: Summer 2005) CauchyEuler En C.F. = A cos X + B sinx COS X, pages denency soln is y = yet YP Cle 3x trexo + 9x2 +24x-2Pages yllty = tanec D (D 2+2) = tan? auxiliary ean is m2+1=0 m = ti Yc- CI CO SOC+02 SIDY y = cosx 42 SIC 41. 421-ylly2 = CosCOS X, P.L. pade Page No. Data I-22 s! t 1 t -t at toot -t2 2 U = loa cosa cos296 2 V- Yltanza v = 8 cosa sin2x dz cos2x SN24osec dr - sin page 7 Data Y = uyl tvyz Y = cosx Clog CcOS) 2 it sing Cloa (secat tana) - sino) y = yet yo y = cl cosx tcasinshere - 32. cosec x de U= page 8 (Page to de 3 Х A-315 4 2 - 4x +2y=0 421= 29 22 Ey 2 = 1, y = 2 Vi= 1 For d₂ = -2 (A-2I) x =pade Page No Date 6.6 know by theo Suppose A has a pair of real eigenvallies di idz and associated eigenvectors vi then gener